How many molecules are in 3.5918 g of sulfuric acid, H2SO4? Show full work. Include units beside every number where applicable. Round your final answer to correct significant figures.



1
Expert's answer
2022-04-26T03:47:04-0400

Solution:

The molar mass of sulfuric acid (H2SO4) is 98.079 g/mol

Therefore,

Moles of H2SO4 = (3.5918 g H2SO4) × (1 mol H2SO4 / 98.079 g H2SO4) = 0.03662 mol H2SO4


One mole of any substance contains 6.022×1023 atoms/molecules.

Therefore,

No of H2SO4 molecules = (0.03662 mol) × (6.022×1023 molecules / 1 mol) = 2.2053×1022 molecules

Number of H2SO4 molecules = 2.2053×1022 molecules


Answer: 2.2053×1022 molecules of sulfuric acid (H2SO4)